OpenStudy (anonymous):

Evaluate. cube root of 7 multiplied by the square root of 7 over sixth root of 7 to the power of 5

OpenStudy (mrnood):

\[\frac{ \sqrt[3]{7} \times \sqrt{7}} { \sqrt[6]{7}^{5} }\] cube root 7 = 7^1/3 sqrt 7 = 7^1/2 (5sixth root 7 ) = 7 ^5/6 use addition and subtraction of indices to simplify

OpenStudy (mrnood):

correction: ^^ (sixth root 7)^5 = 7^5/6